Grade 12th passMechanics

Consider a planet in some solar system which has a mass double the mass of the earth and density equal to the average density of the earth. An object weighing W on earth will weigh a)W b)2W c) W/2 d) 2^1/3 W.....at the planet

Aditya
10 Years agoGrade 12th pass

1 Answer

Riddhish Bhalodia
10 Years ago
W = GmM/R^2 (the values on earth)

now on the planet
M_p = 2M
and density is same hence
M_p/(R_p)^3 = M/R^3
R_p = (cube root 2) * R

Thus the
W_p = [2/(cube root 4)]*W = (cuberoot 2)W => option d